clayton is 4 years younger than his brother josh. if the sum of their ages is 28, how old is each boy?
Answer by addingup(3677)

J = C+4
J+C = 28 substitute for J:
C+4+C = 28
2C = 24
C = 12 Clayton is 12, and his brother Josh is 12+4=16